# One missing piece in Vision and Language: A Survey on Comics Understanding

### v1

@@ 10.2.2 Video generation
Recent advancements have seen significant progress in generating videos from both images [224, 225] and text [226], although accurately defining metrics and benchmarks remains challenging [79]. These developments are particularly notable in the broader Vision and Language community, showcasing the fast-paced evolution of the field.
The domain of comics, however, is still in the early stages of exploring video generation capabilities. In one of the initial attempts, Cao et al. [227] introduced a framework to animate manga panels. This method involves classifying panels based on motion and emotion and using these classifications to animate backgrounds and characters. The final video creation considers both intra-panel animations and transitions between panels. Later, Gupta et al. [228] further advanced this concept with the Comic-to-Video Network (C2VNet). This framework creates panel-audio videos by cropping comic panels and adding a highlighting effect to speech balloons, simulating characters’ dialogues. Alongside this, they introduced the IMCDB, an annotated dataset of Indian Mythological Comics in English, facilitating research in this area.
While these efforts mark initial steps, they still fall short of the more advanced capabilities seen in general Vision-Language video generation, such as producing fully animated video clips with sound from a single image. The potential for integrating comics with video and audio elements presents an exciting opportunity for enhancing storytelling and audience engagement in the comic medium.
Future research in this area could lead to new forms of interactive and immersive comic experiences, blending traditional art with contemporary digital technologies.
10.2.3 Narrative-based Comic Generation
@@
The pioneering efforts in the realm of narrative-based comic generation commenced with the introduction of StoryGAN by Li et al. [229], which marked a significant advancement in generating image sequences from storylines using a sequential conditional GAN framework. This innovation incorporated a Context Encoder, which dynamically tracked the story flow, laying the groundwork for future developments in this area.
Building on this foundation, Maharana et al. [230] enhanced visual quality and coherence by incorporating a dual learning framework that utilized video captioning, a copy-transform mechanism, and MART-based transformers. This approach significantly improved the semantic alignment between the story and generated images, thus enriching the story visualization process.
In terms of character coherence and continuity, Chen et al. [231] focused on maintaining consistent character portrayal throughout the narrative. They adapted Vector-Quantized Variational Autoencoders (VQ-VAE) with a text-to-visual-token architecture, ensuring character coherence in the visual narratives. Furthering this direction, Maharana et al. [211] introduced the concept of story continuation, where the generated visual story is conditioned on a source image, allowing narratives to incorporate new characters more fluidly. They also introduced the DiDeMoSV dataset, which provided a platform for exploring these complex narrative structures.
Recent advancements in story visualization are exemplified by Pan et al. [232], with presenting AR-LDM, a latent diffusion model that significantly raised the standards for visual quality in natural image datasets like VIST. This model was auto-regressively conditioned on history captions and generated images, capturing complex interactions between frames. Rahman et al. [233] introduced a framework with a visual memory module that implicitly captured actor and background context, generating frames that were not only of high visual quality but also consistent with the story. Peng et al. [234] proposed PCSG, a diffusion-based text-to-image synthesis framework with controllable plugins for character consistency, scene layout specification, and character pose specification, further enhancing the personalized aspects of story visualization.
These emerging trends and future directions in narrative-based comic generation showcase a notable shift towards more sophisticated, context-aware, and visually coherent storytelling techniques. The integration of elements like character consistency, scene dynamics, and actor-background interaction is pivotal in creating compelling visual narratives. Future research in this area can further refine these techniques, potentially leading to automated comic creation tools that adapt to various narrative styles and complexities, opening new horizons in the field of comics and visual storytelling.
10.3 Future Tasks
10.3.1 Comics to Scene graph
@@
The last work that tackles this task is [41], which employs a Relationformer architecture [145], explicitly designed for image-to-graph tasks. However, the authors convert the full page into different types of nodes (panels, text, characters) and only consider two types of arches: characters-characters (re-identification) and text-characters (speaker identification). The inherently complexity of the comics, such as the elements in the image (e.g. the background and foreground objects) have not been considered yet. The task has yet to be explored to fine-grained graph creation from single panels, full pages, or a book.
11 CONCLUSION
@@
In reviewing the diverse and innovative efforts in the realm of comic research and technology, it becomes evident that this field is in a state of dynamic and exciting evolution. The array of tasks and applications that researchers have embarked upon reflects not only the complexity of comics as a medium but also their immense potential as a bridge between artistic expression and technological advancement.
We have covered numerous works and categorized them into distinct layers based on a proposed taxonomy. We have aimed to distill the key topics, methods, challenges, and emerging directions that can shape future research in Comics Understanding.
In the future, we will continue to monitor advancements in this field, and we will update our findings in the following repository: https://github.com/emanuelevivoli/awesome-comics-understanding.
A. Methodology details
In this section, we will explore the process of finding the state-of-the-art in Comics and Manga. We utilized the software “Publish or Perish” [^13] Available at: https://harzing.com/resources/publish-or-perish, which comprehends the following Search sources: “Semantic Scholar”, “Google Scholar”, “Web of Science” and “Scopus”. As every source possesses its own interface and query structure, we designed a “SQL-like” query:
  • Title: “comic*”, “manga”
  • Keywords: (“comic*” OR “manga”) AND (“machine*” OR “learning*” OR “ml” OR “computer*” OR “science*” OR “artificial*”)
For every source, we limited the results to the last 1000 from 2010 to 2024. For every paper (out of the 3500 obtained), we used the “Semantic Scholar API” [^14] API at link: https://www.semanticscholar.org/product/api to query the title and obtain the unique DOI. We finally filtered unique papers, automatically removing duplicated copies (DOI) copies, obtaining 2276 papers. This duplicate-removal process has been further implemented also in “zotero” [^15] Zotero software available at: https://forums.zotero.org, fetching data from the websites and using javascript code [^16] Zotero deduplication script. Finally, we define our inclusion/exclusion criteria as follows:
Inclusion. The study regards:
  • (I1)
    Comics and Manga;
  • (I2)
    Computer science;
  • (I3)
    Deep Learning;
  • (I4)
    new Dataset and/or Task.
Exclusion. The study is:
  • (E1)
    not written in English;
  • (E2)
    not available;
  • (E3)
    a duplicate or extension of an already included study;
  • (E4)
    a secondary or tertiary study;
  • (E5)
    in the form of editorials, tutorials, books, extended abstracts, etc.;
  • (E6)
    a non-scientific publication or grey literature;
  • (E7)
    cited at least 2 times (if previous 2022);
  • (E7)
    not published in AI conferences.
After this process, we ended with 216 papers. The last phase of “snowballing” is applied at this point: whi le reading and analyzing the papers we discovered important references that were not in the collection because they were not selected by our query or were discarded by I/E criteria. However, this process is needed to include important papers that are highly cited from our collection. This step enlarged the collection to the final number of 353 papers. Finally, we manually check all the references and include the one that suits the best our taxonomy and scope.
B. Taxonomy
Task classification within the LoCU is informed by both the modalities and dimensions of task inputs and outputs. We define modalities as Images (ranging from no image to multiple pages), Text (from no text to complex multi-sentence structures), Audio (from no audio to complex sound mixes), and Geometry (encompassing points, lines, bounding boxes, segmentations, and 3D geometries). Dimensional analysis considers Space (from non-spatial to tridimensional) and Time (whether time is a factor). Each task is scored based on its input and output characteristics in these modalities and dimensions (see Tables IV and V), with the cumulative score determining its placement within the LoCU layers.
TABLE IV: Modality Scoring for LoCU Tasks
Modality Category Score
Images No image 0
Single panel 1
Single page (multiple panels) 2
Multiple pages 3
Text No text 0
Single word (or tag) 1
Single sentence 2
Multi sentences 3
Instruct multi-sentences 4
Complex multi-sentence 5
Audio No audio 0
Single sound 1
Multi-elements sound 2
Complex mix of sounds 4
Geometry No 0
Point 1
Line 2
Bounding box 3
Instance segmentation 4
Semantic segmentation (multiple masks) 5
3D object geometry 6
TABLE V: Dimensions Scoring for LoCU Tasks
Dimensions Category Score
Space No 0
2D 1
3D 2
Time Not considered 0
Considered 1
In Table VI are reported the values of modality and dimensions, together with the Layer of Comics Understanding Table I. This shows the modalities and dimensions value assigned to every task’s input and output, in order to obtain an objective complexity score that could sort each of the tasks. The acronyms “I, T, A, G” represent “Images, Text, Audio and Geometry”, respectively. Similarly, the acronyms “S, T” correspond to the Dimensions “Space, and Time”, respectively.
TABLE VI: Detailed view of Modalities and Dimensions for both inputs and outputs.
Mods Dim Mods Dim
Layer Category Tasks Input I T A G S T Output I T A G S T Score
0 View Basic Image Viewing (BIV) Text Command 0 0 0 0 0 0 Image Display 0 0 0 0 0 0 0
Image Classification (I-CLS) Image 1 0 0 0 0 0 Tag 0 1 0 0 0 0 2
Emotion Classification Comic Panels/Images 1 0 0 0 0 0 Emotion Labels 0 1 0 0 0 0 2
Action Detection Multiple Panels 2 0 0 0 1 1 Tag 0 1 0 0 0 0 5
Tagging Page Stream Segmentation (PSS) Images 2 0 0 0 1 0 Tags sequence 0 1 0 0 0 0 4
Image Super-Resolution (ISR) Image 1 0 0 0 1 0 Image 1 0 0 0 1 0 4
Style Transfer (ST) Image 1 0 0 0 1 0 Image 1 0 0 0 1 0 4
Vectorization Comic Panels/Images 1 0 0 0 1 0 Vector image 1 0 0 1 1 0 5
1 Augmentation Depth Estimation Comic Panels/Images 1 0 0 0 1 0 Depth Map 1 0 0 0 1 0 4
Object Detection Tag/s + Image 1 1 0 0 1 0 Bounding Boxes 0 0 0 2 1 0 6
Character Re-identification (Character ID) Multiple Panels 2 1 0 0 1 0 Text [Tagging] 0 1 0 2 0 0 7
Grounding Grounding (IG) [prompt] + Image 1 2 0 0 1 0 Bounding Boxes 0 0 0 2 1 0 7
Character-Balloon Association (Speaker ID) Character + Balloons 2 0 0 2 1 0 Text [Tagging] 0 0 0 2 0 0 7
Dialog transcription Image 1 0 0 0 1 1 Text 0 3 0 0 0 1 7
Analysis Translation Image 1 3 0 0 1 0 Text 0 3 0 0 0 0 8
2 Segmentation Instance Segmentation (IS) [prompt] + Image 1 0 0 2 1 0 Segments 0 0 0 3 1 0 8
Image-Text Retrieval (IR) Text 1 3 0 0 1 0 Image 0 2 0 0 1 0 8
Text-Image Retrieval (TR) Image 3 1 0 0 1 0 Text 2 0 0 0 1 0 8
Retrieval Composed Image Retrieval (CIR) Text + Image 3 2 0 0 1 0 Image 2 0 0 0 0 0 8
Image Inpainting (II) Text + [prompt] + Image 1 2 0 2 1 0 Image 2 0 0 0 1 0 9
3 Modification Image Editing via Text (IET) Text + Image 1 4 0 0 1 0 Image 2 0 0 0 1 0 9
Visual Dialog (VisDial) Image + Dialog + Text 2 4 0 0 1 1 Text 0 3 0 0 1 1 13
Visual Reasoning (VR) Image + Text 2 3 0 0 1 1 Text 0 3 0 0 1 1 12
4 Understanding Visual Question Answering (VQA) Image + Text 2 3 0 0 1 1 Text 0 2 0 0 1 0 10
Image Captioning [img-2-text] (IC) Image 1 0 0 0 2 0 Text 0 3 0 0 2 0 8
Image Generation [text-2-img] (IG) Text 0 5 0 0 1 0 Image 1 0 0 0 2 0 9
Scene Graph Generation for Captioning Comic Panel 1 0 0 0 2 0 Scene Graph 0 3 0 3 2 0 11
Generation Sound Generation from Single Panel Single Comic Panel 1 0 0 0 2 0 Sound/Audio 0 0 4 0 3 1 11
3D Model Generation from Images (3DGI) Collection of Images 3 0 0 0 1 0 3D Model 0 0 0 3 3 0 10
Video Generation from Text (VGT) Complex Long Text 0 5 0 0 1 1 Video 2 1 0 0 1 1 12
5 Synthesis Narrative-Based Complex Scene Generation (NCSG) Detailed Narrative Text 2 5 0 0 1 1 Series of Images 3 1 0 0 1 1 15
[Uncaptioned image] Emanuele Vivoli is pursuing its joint-PhD at the Computer Vision Center (Autonomous University of Barcelona) and at the MICC Lab (University of Florence). His research interests encompass multi-modal learning (vision and language integration), with a focus on comics. His work covers a diverse range of topics, including document and table analysis, understanding comics and manga, and creating datasets and frameworks. If you are interested, read his papers.
[Uncaptioned image] Andrey Barsky is a postdoctoral researcher at the Computer Vision Center, Barcelona, Spain. He received his Ph.D. in 2015 from the University of Nottingham in the UK. His research focuses on computer vision and multimodal learning, as well as robustness and explainability in AI models.
[Uncaptioned image] Mohamed Ali Souibgui is a postdoctoral researcher at Computer Vision Center, Barcelona, Spain. He received the Ph.D. degree in 2022 from the Universitat Autònoma de Barcelona (UAB), Spain. His current research focuses on document image analysis using computer vision and machine learning tools, with the focus on privacy and explainability. He is an author of several journal and peer-reviewed conference papers.
[Uncaptioned image] Artemis Llabrés is a PhD student in the Computer Vision Center. She is part of the Vision, Language, and reading group where she is working on multimodal models for document understanding. In 2024 she was awarded best student paper by the 16th IAPR International Workshop on Document Analysis Systems, for her work titled “Image-text matching for large-scale book collections”.
[Uncaptioned image] Marco Bertini is an Associate Professor in Computer Science at the University of Florence, Italy. He is the director of the Media Integration and Communication Center of the University of Florence. His interests are focused on multimedia and computer vision. On these subjects he has addressed semantic analysis, automatic content indexing, semantic retrieval and video quality improvement, applying these techniques to different domains among which cultural heritage. He is author of more than 30 journal papers and more than 140 peer-reviewed conference papers. He has been involved in 10 EU research projects as WP coordinator and researcher, among which IM3I, euTV, ORUSSI, UMETECH, AI4Media and ReInHerit. He has been general and program co-chair of several conferences on multimedia. He is co-founder of Small Pixels, an academic spin-off working on GenAI solutions to improve video quality and video compression.
[Uncaptioned image] Dimosthenis Karatzas is a professor at the Universitat Autónoma de Barcelona and associate director of the Computer Vision Centre, leading the Vision and Language research group. His research focuses on computer vision and machine learning, with an emphasis on robust reading systems and document image analysis. He received the 2013 IAPR/ICDAR Young Investigator Award. He has created two spin-off companies. He chaired the IAPR TC11 on Reading Systems from 2016 to 2021, he is an IEEE senior member, an ELLIS fellow and co-director of the ELLIS Unit Barcelona, and a member of the Artificial Intelligence Doctoral Academy (AIDA) Research and Industry Board. He launched the Robust Reading Competition portal, which serves more than 50,000 researchers worldwide.

### v2

@@ 10.2.2. Video generation
Recent advancements in video generation from both images (Yan et al., 2023; Wang et al., 2023a) and text (Gupta et al., 2023) have made significant strides, though defining accurate metrics and benchmarks remains a challenge (Fan et al., 2024). These developments are rapidly evolving within the broader Vision and Language community. However, in the domain of comics, video generation is still in its early stages. One of the first attempts came from Cao et al. (Cao et al., 2017), who introduced a framework to animate manga panels. Their method classifies panels based on motion and emotion, using this information to animate characters and backgrounds. The resulting video includes both intra-panel animations and transitions between panels. Later, Gupta et al. (Gupta et al., 2021) expanded on this with C2VNet, which generates panel-audio videos by cropping panels and adding a highlighting effect to speech balloons to simulate character dialogue. They also introduced IMCDB, a dataset of annotated Indian Mythological Comics in English, to support research in this field. While these efforts represent initial steps, they are still far from the capabilities seen in Vision-Language video generation, such as producing fully animated video clips with sound from a single image. The integration of comics with video and audio elements offers exciting potential for enhancing storytelling and audience engagement. Future research could pave the way for interactive and immersive comic experiences, blending traditional art with modern digital technologies.
10.2.3. Narrative-based Comic Generation
@@
The pioneering efforts in the realm of narrative-based comic generation commenced with the introduction of StoryGAN by Li et al. (Li et al., 2018b), which marked a significant advancement in generating image sequences from storylines using a sequential conditional GAN framework. This innovation incorporated a Context Encoder, which dynamically tracked the story flow, laying the groundwork for future developments in this area. Building on this foundation, Maharana et al. (Maharana et al., 2021) enhanced visual quality and coherence by incorporating a dual learning framework that utilized video captioning, a copy-transform mechanism, and MART-based transformers. This approach significantly improved the semantic alignment between the story and generated images, thus enriching the story visualization process. In terms of character coherence and continuity, Chen et al. (Chen et al., 2022) focused on maintaining consistent character portrayal throughout the narrative. They adapted Vector-Quantized Variational Autoencoders (VQ-VAE) with a text-to-visual-token architecture, ensuring character coherence in the visual narratives. Furthering this direction, Maharana et al. (Maharana et al., 2022) introduced the concept of story continuation, where the generated visual story is conditioned on a source image, allowing narratives to incorporate new characters more fluidly. They also introduced the DiDeMoSV dataset, which provided a platform for exploring these complex narrative structures. Recent advancements in story visualization are exemplified by Pan et al. (Pan et al., 2022), with presenting AR-LDM, a latent diffusion model that significantly raised the standards for visual quality in natural image datasets like VIST. This model was auto-regressively conditioned on history captions and generated images, capturing complex interactions between frames. Rahman et al. (Rahman et al., 2022) introduced a framework with a visual memory module that implicitly captured actor and background context, generating frames that were not only of high visual quality but also consistent with the story. Peng et al. (Peng et al., 2023) proposed PCSG, a diffusion-based text-to-image synthesis framework with controllable plugins for character consistency, scene layout specification, and character pose specification, further enhancing the personalized aspects of story visualization. These emerging trends and future directions in narrative-based comic generation showcase a notable shift towards more sophisticated, context-aware, and visually coherent storytelling techniques, potentially leading to automated comic creation tools that adapt to various narrative styles and complexities.
10.3. Future Tasks
10.3.1. Comics to Scene graph
@@
The last work that tackles this task is (Sachdeva and Zisserman, 2024), which employs a Relationformer architecture (Shit et al., 2022), explicitly designed for image-to-graph tasks. However, the authors convert the full page into different types of nodes (panels, text, characters) and only consider two types of arches: characters-characters (re-identification) and text-characters (speaker identification). The inherently complexity of the comics, such as the elements in the image (e.g. the background and foreground objects) have not been considered yet. The task has yet to be explored to fine-grained graph creation from single panels, full pages, or a book.
11. CONCLUSION
@@
In reviewing the diverse and innovative efforts in the realm of comic research and technology, it becomes evident that this field is in a state of dynamic and exciting evolution. The array of tasks and applications that researchers have embarked upon reflects not only the complexity of comics as a medium but also their immense potential as a bridge between artistic expression and technological advancement. We have covered numerous works and categorized them into distinct layers based on a proposed taxonomy. We have aimed to distill the key topics, methods, challenges, and emerging directions that can shape future research in Comics Understanding. In the future, we will continue to monitor advancements in this field, and we will update our findings in the following repository: https://github.com/emanuelevivoli/awesome-comics-understanding.